OpenStudy (anonymous):

1/3(-1)+1 1/3(0)+1 1/3(1)+1 1/3(2)+1

jhonyy9 (jhonyy9):

-1/3 +1 =(-1+3)/3=2/3 0/3 +1 =0+1=1 1/3 +1 =(1+3)/3 = 4/3 2/3 +1 =(2+3)/3 = 5/3